OpenStudy (anonymous):

here A solution to the compound inequality 16 > 4 - 3x > -11 is provided below. Choose the statement that is true about the given solution. 16 > 4 - 3x and 4 - 3x > -11 Line 1 12 > -3x and -3x > -15 Line 2 -4 > x and x > 5 Line 3

OpenStudy (anonymous):

@Shadowys @mayankdevnani

OpenStudy (anonymous):

some one help

OpenStudy (anonymous):

here is a mistake in line # 1. There is a mistake in line # 2. There is a mistake in line # 3. There are no mistakes, the solution is correct.

OpenStudy (anonymous):

thoses are the choises

OpenStudy (anonymous):

mist in line 3

OpenStudy (anonymous):

ok thanks you help me with some more plz

OpenStudy (anonymous):

Well, first we view the first line. 16 > 4 - 3x and 4 - 3x > -11 Since the word is 'and', and that means it can be combined to one inequality. 16 > 4 - 3x > -11. So it's right. Line two. 12 > -3x and -3x > -15 this is done by subtracting four from the first line. Again, no mistake. Line three -4 > x and x > 5 Now, the x > 5 is wrong as you'll need to reverse the signs when you divide by a negative number. So line three is wrong.

OpenStudy (anonymous):

thanks!!!!!!!!!!!!! shadowwyyy In Lake Kerr, a black bass must be less than or equal to 15 inches, or longer than or equal to 24 inches for the fish to be considered a keeper. Choose the combined inequality that describes this scenario. 15 less than or greater to x less than or greater to 24 x less than or greater to 15 or x greater than or equal to 24 x less than or greater to 24 none of the above

OpenStudy (anonymous):

As described, we want, 15<=x<=24 None of the above describes such a scenario. It has no equal to...

OpenStudy (anonymous):

If the greater to word is replaced by equal to, then the answer is the first one.

OpenStudy (anonymous):

another one plzz Solve: 5(x + 6) less than or greater to 3x - 4 x greater than or equal to -17 x greater than or equal to 17 x less than or greater to -17 x less than or greater to -5

OpenStudy (anonymous):

5(x+6)<=3x-4 5x+30<=3x-4 2x<=-34 x<=-17 So, what do you think ? :)

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

Yup! :) You've got it.

OpenStudy (anonymous):

and one more plzz if you have time

OpenStudy (anonymous):

Solve -4x - 5 > -9 and describe the graph of the solution. x > 1; closed circle on 1, shading to the right x < 1; closed circle on 1, shading to the left x < 1; open circle on 1, shading to the left x > 1; open circle on 1, shading to the right

OpenStudy (anonymous):

What did you get after solving it? x<1?

OpenStudy (anonymous):

x < 1; open circle on 1, shading to the left

OpenStudy (anonymous):

ya what he said lol hahahah

OpenStudy (anonymous):

Lol it's right.

OpenStudy (anonymous):

thank you thank you thank you

OpenStudy (anonymous):

i like you hahahaha

OpenStudy (anonymous):

-4y + 17 less than or greater to 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution. i dont understand this one either

OpenStudy (anonymous):

rewrite the question plz its less than or greater than???

OpenStudy (anonymous):

-4y + 17 <=29 y>=-3 So the graph is  closed circle , shaded top.

OpenStudy (anonymous):

its a sing going this way < with a line under it

OpenStudy (anonymous):

so @Shadowys it would be a closed circle corect

OpenStudy (anonymous):

yup! :) it's a \(\le\) which is lesser or equal to

OpenStudy (anonymous):